# Learning to Route for Dynamic Adapter Composition in Continual Learning with Language Models

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) methods are increasingly used with pre-trained language models (PLMs) for continual learning (CL). These methods involve training a PEFT module for each new task and using similarity-based selection to route modules during inference. However, they face two major limitations: 1) interference with already learned modules and 2) suboptimal routing when composing modules. In this paper, we introduce a method that isolates the training of PEFT modules for task specialization. Then, before evaluation, it learns to compose the previously learned modules by training a router that leverages samples from a small memory. We evaluate our method in two CL setups using several benchmarks. Our results show that our method provides a better composition of PEFT modules, leading to better generalization and performance compared to previous methods.
@@ 1 Introduction
PEFT methods, like prompts or adapters, are lightweight modules that can be tuned for downstream tasks while keeping the PLM frozen. Recent work shows that these methods can be competitive with, or even superior to, full fine-tuning of PLMs Li and Liang (2021); Hu et al. (2022). Additionally, research Wang et al. (2023, 2024, 2022b); Smith et al. (2023) demonstrates their effectiveness in CL.
However, PEFT approaches have two significant limitations: 1) When a new task arises, a new PEFT module is added, and the previously learned ones remain frozen but activated Razdaibiedina et al. (2023). Previous modules may interfere with the optimal task knowledge acquisition by the current PEFT module. 2) They rely on a similarity-based selection mechanism Wang et al. (2024) that may not accurately represent a routing function that effectively composes the PEFT modules.
To address these issues, we introduce Learning to Route for dynamic PEFT composition (L2R). L2R is a simple but effective method that trains PEFT modules in isolation to avoid interference from previous knowledge when learning new tasks Wang et al. (2023). Then, it uses previous examples stored in memory to learn a routing function for dynamically composing PEFT modules during inference, a process analogous to local adaptation d'Autume et al. (2019). Our extensive evaluations show that L2R is competitive against other PEFT-based methods across benchmarks and CL setups.
@@ 2 Related Work
Recent techniques use PEFT for CL, adding new modules as tasks arise while freezing previous ones to maintain a strong backbone and leverage past knowledge Razdaibiedina et al. (2023). This has been extended for models to learn module composition for inference Wang et al. (2024).
@@ 3 Method
We consider a CL setup for NLP d'Autume et al. (2019), where a model needs to learn from a sequence of $T$ tasks: $D=\{D_{1},...,D_{T}\}$. Each task $t$ consists of a new data distribution $D_{t}=(X_{t},Y_{t})$, where $X_{t}$ are the input instances and $Y_{t}$ are the labels. A model in this setup consists of a PLM that processes input $x$ and a classifier for predicting class $y$. We extend this setup by PEFT modules, specifically adapters, which consistently achieve superior predictive performance than other techniques in CL Wistuba et al. (2024).
@@
L2R employs a set of adapters $A=\{A_{1},...,A_{T}\}$ that learn task-specific knowledge from a data stream (Figure 1a). Specifically, an adapter $A_{t}$ is activated and trained when a new task $D_{t}$ emerges. Previous adapters are deactivated, allowing $A_{t}$ to learn in isolation and specialize in task $t$ Wang et al. (2023). We keep the backbone frozen and train only the adapter $A_{t}$. For simplicity, we omit the fact that each task uses $L$ adapters that match the number of PLM layers.
Adapters offer significant benefits. Their robust modeling capabilities allow them to effectively capture distribution-specific details from each task and the high parameter efficiency enables efficient training of our method whenever a new task emerges.
@@
Our method is equipped with a non-parametric memory $M$ that stores previously seen training examples. A similar component is used in replay methods d'Autume et al. (2019), where memory is used to perform replay during training. However, our memory is leveraged prior to test inference to learn a routing function resembling the local adaptation to improve generalization.
We determine the memory capacity as a percentage of the total dataset size and populate it by randomly sampling the training data. Random sampling captures the global distribution of the data stream Araujo et al. (2022a), which will help in the effective training of routers. While more advanced memory population methods could enhance performance Hurtado et al. (2023); Hayes and Kanan (2021), we leave this exploration for future work.
@@
Once the adapters $A$ have been trained, they may be combined for inference (Figure 1b). For this, we add a router network $R$ to each adapter across the model. This router computes the probability of the input being sent to each adapter. Our router takes the [cls] representation as input to decide how to route the whole input. This provides two advantages: 1) the routers learn to route based on input structure, and 2) can produce different combinations at different layers based on the model’s abstraction hierarchy.
Instead of using a softmax across adapters, where modules compete for activation, we adopt an approach inspired by Ponti et al. (2023), emphasizing task-level modularity that respects task hierarchy, where more complex tasks encompass simpler ones as sub-tasks. Each output of the router is a binary scalar indicating whether an adapter is active for a given input. This is implemented as a collection of Bernoulli distributions, relaxed through a Gumbel-sigmoid, which ensures stochasticity.
To obtain a competent router, we propose a memory-based router learning process. Inspired by local adaptation, we leverage the elements in memory to train the router network parameters. However, unlike local adaptation that adapts the model for each input, our method performs this process only once after the training phase to learn the routing functions. After this, the model is ready to perform inference at any time.
@@
Given routing probabilities, we can compose our adapters in different ways. We consider two options: 1) a weighted average of their outputs Shazeer et al. (2017b), denoted as L2R-wavg, and 2) merging adapter parameters via a weighted average of their weights Wortsman et al. (2022), referred to as L2R-merge.
@@ 4 Experimental Setup
Baselines
Implementation Details
Figure 2: L2R-wavg performance across tasks and CL setups. Only order 3 is shown for MTL5 and AfriSenti.
@@ 5 Results
In this section, we introduce our results with a summarized Table 1. For full results, see Appendix C, and for an analysis of efficiency, see Appendix D.
@@
(a) L2R-wavg.
(b) L2R-wavg (Softmax).
Figure 3: Average router scores on task 5 of the MTL5 test set using (a) Gumbel-sigmoid and (b) Softmax.
Router Analysis
As discussed in section 3, we utilize Gumbel-sigmoid activation in our routers to theoretically enhance task-level modularity over activation competency offered by Softmax. To validate this, we compare the routing probabilities from L2R-wavg in CIL with the scores from a version trained using Softmax activation. Figure 3 shows the router probabilities generated for the task 5 test set in MTL5. We observe that L2R-wavg scores activate task 5 moderately, but also leverage knowledge from other tasks. In contrast, the Softmax version predominantly activates the adapter for task 5. These results confirm that L2R provides a composition that leverages knowledge from another task and not only relies on a unique adapter. We found similar behavior across tasks.
@@ 6 Conclusion
We introduce L2R, a method for learning to route in CL that enables a PLM to dynamically combine adapters, providing better generalization and performance. Our experiments across various benchmarks demonstrate the proficiency of our method in TIL and in the challenging CIL setup.
@@ Appendix A Additional Benchmark Details
Web-of-science (WOS) Kowsari et al. (2017) initially began as a hierarchical dataset for categorizing documents. It includes research papers from seven distinct fields: biochemistry, civil engineering, computer science, electrical engineering, medical science, mechanical engineering, and psychology. Each of these fields represents a high-level category for document classification, with multiple subcategories within each field. This dataset has about 11,967 instances and 33 classes. In line with Wang et al. (2023), we split to train/val/test set in the ratio of 0.6:0.2:0.2 and structured 7 sequential learning tasks based on its high-level categories:

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) methods are increasingly used with pre-trained language models (PLMs) for continual learning (CL). These methods typically involve training a PEFT module for each new task and employing similarity-based selection to route modules during inference. However, they face two major limitations: 1) interference during module training with already learned modules and 2) suboptimal routing when composing modules. In this paper, we present L2R, a method that isolates the training of new PEFT modules to ensure their task specialization. L2R then learns to compose the learned modules by training a network of routers that leverages a small memory containing examples of previously seen tasks. We evaluate our method in two CL setups using various benchmarks. Our results demonstrate that L2R provides an effective composition of PEFT modules, leading to improved generalization and performance compared to other methods.
@@ 1 Introduction
PEFT methods, like prompts or adapters, are lightweight modules that can be trained for downstream tasks while keeping the PLM frozen. Recent work shows that these methods can be competitive with, or even superior to, full fine-tuning of PLMs Li and Liang (2021); Hu et al. (2022). Additionally, research Wang et al. (2023, 2024, 2022b); Smith et al. (2023) demonstrates their effectiveness in CL.
However, PEFT approaches for CL have two significant limitations: 1) When a new task arises, a new PEFT module is added, and the previously learned ones remain frozen but activated Razdaibiedina et al. (2023). Previous modules may interfere with optimal task knowledge acquisition by the current PEFT module. 2) They rely on a similarity-based selection mechanism Wang et al. (2024) that may not accurately represent a routing function that effectively composes the PEFT modules.
To address these issues, we introduce Learning to Route for dynamic PEFT composition (L2R). L2R is a simple but effective method that trains PEFT modules in isolation to avoid interference from previous knowledge when learning new tasks Wang et al. (2023). Then, it uses previous examples stored in a memory to learn a routing function for dynamically composing PEFT modules during inference, a process analogous to local adaptation d'Autume et al. (2019). Our extensive evaluations show that L2R is competitive against other PEFT-based methods across benchmarks and CL setups.
@@ 2 Related Work
Recent techniques use PEFT for CL, adding new modules as tasks arise while freezing previous ones to maintain a robust backbone and leverage past knowledge Razdaibiedina et al. (2023). This approach has been extended to learn the composition of PEFT modules based on similarity Wang et al. (2024). While effective, these methods may suffer from interference during learning new modules and suboptimal module composition during inference.
@@ 3 Method
We consider a CL setup for NLP d'Autume et al. (2019), where a model needs to learn from a sequence of $T$ tasks: $D=\{D_{1},...,D_{T}\}$. Each task $t$ consists of a new data distribution $D_{t}=(X_{t},Y_{t})$, where $X_{t}$ are the input instances and $Y_{t}$ are the labels. A model in this setup consists of a PLM that processes input $x$ and a classifier for predicting class $y$. We extend this setup with PEFT modules, specifically adapters, which consistently achieve superior predictive performance than other techniques in CL Wistuba et al. (2024).
@@
L2R employs a set of adapters $A=\{A_{1},...,A_{T}\}$ that learn task-specific knowledge from a data stream (Figure 1a). Specifically, an adapter $A_{t}$ is activated and trained when a new task $D_{t}$ emerges. Previous adapters are deactivated, allowing $A_{t}$ to learn in isolation and specialize in task $t$ Wang et al. (2023). We keep the backbone frozen and train only the adapter $A_{t}$ on the current task $t$. To simplify the explanation, we omit the fact that each task uses $L$ adapters that match the number of PLM layers.
Adapters offer significant benefits Wistuba et al. (2024). Their robust modeling capabilities allow them to effectively capture distribution-specific details from each task and the high parameter efficiency enables efficient training of our method whenever a new task emerges.
@@
Our method is equipped with a non-parametric memory $M=\{(x_{i},y_{i})\}$ that stores input-output pairs of previously seen training examples. A similar component is used in replay methods d'Autume et al. (2019), where memory is used to perform replay during training. However, our memory is leveraged prior to test inference to learn a routing function resembling the local adaptation process to improve generalization.
We determine the memory capacity as a percentage of the total dataset size and populate it by randomly sampling examples from the training data $M=\{\texttt{sample}(D_{1}),...,\texttt{sample}(D_{T})\}$. Random sampling effectively captures the global distribution of the data stream and has proven successful in experience replay Araujo et al. (2022a), so we hypothesize that it will similarly support router training. While more advanced memory population methods could enhance router performance Hurtado et al. (2023); Hayes and Kanan (2021), we leave this exploration for future work.
@@
Once the adapters $A$ have been trained, they can be combined for inference (Figure 1b). For this, we implement a router network $R$ within the model. Note that, for simplicity, we omit the detail that a separate $R$ is added for each of the $L$ layers. This router computes the probability of the input being sent to each adapter. This provides two advantages: 1) the routers learn to route based on input structure, and 2) can produce different combinations at different layers based on the model’s abstraction hierarchy.
In practice, the router takes the [cls] representation as input to generate an allocation vector $z=R(h_{[\textsc{cls}]})$, which is then used to compute the output distribution. Rather than applying Softmax across adapters, where modules compete for activation, we adopt an approach inspired by Ponti et al. (2023), emphasizing task-level modularity that respects task hierarchy, with more complex tasks encompassing simpler ones as sub-tasks. In this case, the router would output a binary scalar for each adapter, indicating whether it is active for a given input. As a binary vector is not differentiable, this is implemented as a collection of Bernoulli distributions, relaxed through a Gumbel-sigmoid to ensure stochasticity: $\hat{z}_{t}=\sigma\log\left[\frac{\sigma(z_{t})u}{(1-\sigma(z_{t}))(1-u)}\right]$, where $u\sim\texttt{Uniform}(0,1)$.
To obtain a competent router, we propose a memory-based router learning process. Inspired by local adaptation, we leverage the elements in memory to train the router network parameters. However, unlike local adaptation, which adjusts the model for every inference example, our method performs this process only once after the training phase to learn the routing functions, enabling inference on any example without further adaptation.
@@
Given routing probabilities, the model can compose its adapters in different ways. We consider two options: 1) a weighted average of their outputs Shazeer et al. (2017b), denoted as L2R-wavg, and 2) merging adapter parameters via a weighted average of their weights Wortsman et al. (2022), referred to as L2R-merge.
Note that these approaches differ slightly. The first uses all adapters for inference and returns a weighted average of the hidden state outputs: $H_{wavg}=\sum_{t=1}^{T}\hat{z}_{t}*H_{t}$, where $H_{t}=A_{t}(x)$, while the second merges all adapters into a single one to be used to produce a unique hidden state output: $A_{merge}=\sum_{t=1}^{T}\hat{z}_{t}*A_{t}$.
@@ 4 Experimental Setup
Figure 2: L2R-wavg performance across tasks and CL setups. Results for order 3 are shown for MTL5 and AfriSenti, and for order 1 for WOS.
@@ 5 Results
In this section, we present our results, summarized in Table 1, which reports the average performance across orders for each benchmark. Full results are provided in Appendix C. Additionally, a brief efficiency analysis is available in  D.
@@
Method: L2R-wavg MTL5 WOS AfriSenti
Gumbel-sigmoid 78.0 79.90 60.04
Softmax 77.8 60.74 59.90
Table 2: Accuracy results on MTL5, WOS, and AfriSenti benchmarks for the CIL setup of L2R-wavg using Gumbel-sigmoid and Softmax.
Figure 3: Average router scores for task 2 of MTL5 (order 4), task 3 of WOS (order 1), and task 4 of AfriSenti (order 1) using Gumbel-sigmoid (top) and Softmax (bottom). Scores were computed on the test sets.
@@
Impact of Gumbel-sigmoid Distribution
As discussed in section 3, we use Gumbel-sigmoid distribution during the training of the routers to theoretically enhance task-level modularity over activation competency offered by Softmax. To validate this, we compare the resulting routing scores from L2R-wavg in CIL with the ones from a version trained using Softmax function.
Table 2 shows the results, highlighting that using Gumbel-sigmoid during the training of the routing function consistently outperforms the use of Softmax activation. The improvement is observed across all benchmarks, with accuracy gains ranging from 0.1 to 19.2 points. The WOS benchmark exhibits the most significant improvement, likely because it consists of more tasks, which allows for more modules to be leveraged.
Figure 3 shows router probabilities for the test sets of MTL5 (order 4), WoS (order 1), and AfriSenti (order 1). On the left, for MTL5 task 2 (news classification), adapter 2 is the most activated in both model versions. However, the Softmax version predominantly uses adapter 2, while the Gumbel-sigmoid version also leverages adapters 1 and 4, which specialize in sentiment analysis.
In the middle, for the WOS task 3, both models activate multiple adapters because all tasks are document categorization across various domains. With Gumbel-sigmoid, adapter 3 (the correct one) is primarily activated, while the Softmax version strongly activates adapter 7, with less use of others.
On the right, for AfriSenti task 4 (sentiment analysis with varying input languages), the Gumbel-sigmoid version strongly activates adapters 3 and 4, with lighter activation of the rest. Adapters 3 and 4 correspond to Hausa and Igbo, both spoken in Nigeria, explaining their activation. Adapter 4 is slightly more activated, showing the router’s ability to detect the target task while using other adapters. In contrast, the Softmax version primarily activates adapter 3, with less activation of others.
@@ 6 Conclusion
We introduced L2R, a method for routing in CL that enables PLMs to dynamically combine adapters, improving generalization and performance. Experiments across benchmarks demonstrate L2R’s effectiveness in both TIL and the challenging CIL setting. We also show that larger memory enhances routing function learning and that L2R effectively leverages diverse adapter combinations.
@@
7 Acknowledgements
This work was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme: CALCULUS project (Grant Agreement No. 788506) and KeepOnLearning project (Grant Agreement No. 101021347).
@@ Appendix A Additional Benchmark Details
Web-of-science (WOS) Kowsari et al. (2017) initially began as a hierarchical dataset for categorizing documents. It includes research papers from seven distinct fields: biochemistry, civil engineering, computer science, electrical engineering, medical science, mechanical engineering, and psychology. Each of these fields represents a high-level category for document classification, with multiple subcategories within each field. This dataset has about 11,967 instances and 33 classes. In line with Wang et al. (2023), we split to train/val/test set in the ratio of 0.6:0.2:0.2 and structured 7 sequential learning tasks based on the high-level categories of the dataset: